anne and nancy use a metal alloy that is 17 percent copper to make jewerly. how many ounces of a 15 percent alloy must be mixed with a 18 percent alloy to form 78 ounces of the desired alloy?
Answer by TimothyLamb(4379)

x = ounces of 15% alloy
y = ounces of 18% alloy
---
x + y = 78
15x + 18y = 17*78
---
put the system of linear equations into standard form
---
x + y = 78
15x + 18y = 1326
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = ounces of 15% alloy = 26
y = ounces of 18% alloy = 52
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---